# One Language, Many Gaps: Evaluating Dialect Fairness and Robustness of Large Language Models in Reasoning Tasks

### v1

@@ 3.1 Experimental Setting
We test four families of models, one proprietary and three open-source, on zero-shot prompting and zero-shot Chain of Thought (Kojima et al., 2022; Wei et al., 2022) to simulate one setting for general users and one setting for expert users. We deliberately do not test more advanced prompting methods such as Tree of Thought (Yao et al., 2024) and Self-Refine (Madaan et al., 2024) as we are interested in how LLMs behave when prompted for daily usage since this is the context in which input is most likely to contain dialectal features.
We elaborate further on model choices in Section 3.1.1 and experiment settings in Section 3.1.2.
@@ 3.1.1 Models
Here, we report the details about the models we test. The rationale is to benchmark widely used LLMs with impressive reasoning performance.
@@
We use GPT-4o, GPT-4, GPT-3.5-turbo (Achiam et al., 2023), [^3] https://openai.com/index/hello-gpt-4o/, https://openai.com/index/gpt-4/,    https://platform.openai.com/docs/models/gpt-3-5-turbo. as a family of SotA closed-source models to compare with open-source models for dialect robustness.
@@
We use Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). Mistral-7B-Instruct-v0.3 is reported to be outstanding in reasoning; with Mixtral-8x7B-Instruct-v0.1, we can understand whether Mixture-of-Expert architectures enhance dialect robustness.
@@
We use Phi-3-Mini/Small/Medium-128K-Instruct (Abdin et al., 2024; Gunasekar et al., 2023) in our experiment. Phi-3 models, trained on carefully designed “textbook” data, are reported for impressive performance in reasoning despite their small sizes (3.8/7/14B parameters each). We use these models to understand how (i) scaling laws (Kaplan et al., 2020) and (ii) highly curated training data affect LLMs’ dialect robustness and fairness.
@@ 3.1.2 Implementation and Evaluation
We use temperature zero for all experiments to ensure maximum reproducibility. We report two prompting methods in our main results: (i) zero-shot (i.e., directly prompting LLMs with task instances, which resembles general real-life use cases the most) and (ii) zero-shot Chain of Thought (Wei et al., 2022; Kojima et al., 2022) (CoT, i.e., adding instructions in the spirit of “Let’s think step by step” on top of task descriptions, which resembles expert user prompts to improve model performance). [^4] We also test non-zero temperatures and report results in Appendix A.6. We report further implementation details in Appendix A.5.
Model Setting Original AAVE
GPT-4o Zero-shot 0.832 0.716Δ=0.116
CoT 0.826 0.784Δ=0.043
GPT-4 Zero-shot 0.678 0.612Δ=0.067
CoT 0.706 0.590Δ=0.115
GPT-3.5-turbo Zero-shot 0.531 0.460Δ=0.072
CoT 0.517 0.416Δ=0.101
LLaMA-3.1-70B-Instruct Zero-shot 0.663 0.599Δ=0.064
CoT 0.759 0.711Δ=0.049
LLaMA-3-70B-Instruct Zero-shot 0.628 0.562Δ=0.066
CoT 0.693 0.622Δ=0.072
LLaMA-3-8B-Instruct Zero-shot 0.489 0.480Δ=0.009
CoT 0.488 0.472Δ=0.016
Mixtral-8x7B-Instruct-v0.1 Zero-shot 0.388 0.274Δ=0.114
CoT 0.431 0.345Δ=0.086
Mistral-7B-Instruct-v0.3 Zero-shot 0.297 0.214Δ=0.083
CoT 0.305 0.252Δ=0.053
Phi-3-Medium-128K-Instruct Zero-shot 0.513 0.454Δ=0.059
CoT 0.513 0.458Δ=0.055
Phi-3-Small-128K-Instruct Zero-shot 0.530 0.421Δ=0.109
CoT 0.549 0.429Δ=0.119
Phi-3-Mini-128K-Instruct Zero-shot 0.456 0.410Δ=0.046
CoT 0.528 0.461Δ=0.067
Table 2: Pass rates for testing models with zero-shot and CoT prompting on ReDial. We follow the recommendations from Dror et al. (2018) and test the statistical significance of performance differences between Standardized English and AAVE using the McNemar’s test for binary data (McNemar, 1947). We correct p-values for multiple measurements using the Holm-Bonferroni method (Holm, 1979). Results in bold show a statistically significant deviation between AAVE and Standardized ReDial (i.e., models have significant drops in AAVE). We also indicate the absolute delta in performance between the two settings.
Algorithm Math Logic Comprehensive Average
Zero-shot Original 0.602 0.733 0.578 0.191 0.546
AAVE 0.517Δ=0.085 0.665Δ=0.068 0.522Δ=0.056 0.101Δ=0.090 0.473Δ=0.073
CoT Original 0.597 0.811 0.580 0.240 0.574
AAVE 0.495Δ=0.102 0.742Δ=0.068 0.530Δ=0.050 0.177Δ=0.063 0.504Δ=0.070
Table 3: Pass rates by task averaged across responses from all models with zero-shot and CoT prompting. Results in bold show a statistically significant deviation according to McNemar’s tests applied to AAVE and Standardized English (i.e., models have significant drops in AAVE). We also indicate the absolute delta in performance between the two settings.
@@ 3.2 Experimental Results
We report pass rates for ReDial in Table 2 and  3, respectively averaged by task and by model (see detailed results in Appendix A.7). We now summarise the main results of our experiments.
All models are brittle to AAVE.
@@
We find that all models experience performance drops in AAVE compared to Standardized ReDial, and these drops are statistically significant in all cases, with the sole exception of LLaMA-3-8B-Instruct. Except for GPT-4o and LLaMA-3.1-70B-Instruct, all other models have pass rates of similar to or below 0.6 in AAVE Redial, even with CoT, while the best pass rate in Standardized ReDial is 0.832. This indicates that our benchmark poses huge challenges to models, both in terms of absolute performance and with respect to their dialect robustness and fairness.
All reasoning tasks are brittle to AAVE.
LLMs experience the most severe drops in tasks related to algorithm and comprehensive reasoning. In comprehensive tasks that require the composition of more than one elementary reasoning skill, the relative performance drop is especially strong: almost $50\%$ relative performance drop across models with zero-shot, and close to $30\%$ drop with CoT. LLMs face further difficulty when they are asked in a dialect to compose different skills for solving problems.
Scaling does not make models more robust to AAVE.
Comparing within LLaMA-3 and Phi-3 model famlies, we find that although increasing model size improves absolute performance, it cannot close the Standardized English-AAVE performance gaps. For example, comparing LLaMA-3-8B with LLaMA-3-70B, the performance gap in zero-shot widens from $0.009$ to $0.066$ (Table 2) Thus, scaling does not always result in better dialect robustness and fairness. We also find that Mixtral-8x7B-Instruct-v0.1 has an even bigger drop compared to the smaller Mistral-7B-Instruct-v0.3. This suggests that Mixture-of-Experts does not necessarily bring performance gains to models prompted in dialects either.
Highly curated data is particularly brittle to AAVE.
The Phi-3 models, which are trained on highly curated clean data, achieve impressive performance despite their small sizes in Standardized ReDial. For instance, Phi-3-Mini-128K-Instruct (3.8B) outperforms LLaMA-3-8B-Instruct (8B) in the Standardized ReDial with CoT prompting ($0.528$ vs. $0.488$ pass rate). However, it suffers from a large ($0.067$) performance drop in AAVE ReDial, while LLaMA only drops by less than $0.016$, a result that is not statistically significant. This finding is in line with Dodge et al. (2021), which suggests that cleaning and removing data exacerbates unfairness to minority groups.
4 Analysis of Brittleness to AAVE
This section investigates the links between dialectal features and the AAVE brittleness. We compare model performance on human-written AAVE data and misspelled English inputs (Section 4.1) to show that AAVE training data skewness does not explain the whole picture of dialect unfairness and brittleness. We further show that asking a model to rephrase an AAVE input into Standardized English and then answer the question does not cancel the unfairness but tends to increase the computational cost (in terms of tokens generated, Section 4.2). Last, we qualitatively examine cases where LLMs fail in AAVE, even after rephrasing in Standardized English, but succeed in the prompts that are originally written in Standardized English, and identify key error patterns for them (Section 4.3).
4.1 Data Skewness Does not Explain AAVE Brittleness
One possible explanation of the performance drop on AAVE is its infrequency in LLMs’ training corpora. As the model’s training data is largely unknown, we use perplexity as a proxy to measure how familiar the LLMs are with some data: the higher the perplexity, the less familiar an LLM is with the data. We conduct experiments on LLaMA-3.1-70B-Instruct, Phi-3-Medium/Mini-128K-Instruct on Standardized and AAVE ReDial and report their perplexities averaged across instances in Table 4.
LLaMA-3.1-70B-Instruct Phi-3-Medium-128K-Instruct Phi-3-Mini-128K-Instruct
Standardized 9.4 5.9 7.1
AAVE 17.5 12.5 15.9
@@ 4.1 Data Skewness Does not Explain AAVE Brittleness
Table 4: Averaged perplexities across instances calculated by different models on Standardized/AAVE ReDial.
Figure 3: Model performance on misspelled Standardized English compared to human-written AAVE data. We gradually add noise to Standardized ReDial to increase its perplexities until they surpass the perplexity of AAVE ReDial and report the models’ performance on every perturbation level. Horizontal and vertical lines refer to model pass rates/perplexities on AAVE ReDial respectively. Larger LLMs (i.e., LLaMA-3.1-70B-Instruct and Phi-3-Medium-128K-Instruct) perform worse on AAVE than on perturbed text with a similar perplexity level.
As expected, LLMs have higher perplexities on AAVE than Standardized ReDial, which indicates they are indeed less familiar with AAVE than with Standardized English. Does this mean that we can fully attribute the dialect performance gap to its data skewness? To answer this question, we gradually perturb Standardized English by injecting typos, such that we decrease the LLMs’ familiarity with the input texts (i.e., the measured perplexity goes up). Specifically, we simulate typos by replacing/deleting/adding characters in Standardized ReDial. We control an increasing perturbation rate until the tested models’ perplexities exceed those measured in AAVE (i.e., when models are less familiar with misspelled Standardized English than with AAVE). [^5] In practice, we introduce perturbations of densities $\{0,0.02,0.04,0.06\}$, which results in four different typo perplexity levels for each model.
Results are in Figure 3. Interestingly, although LLaMA-3.1-70B-Instruct and Phi-3-Medium-128K-Instruct performance drops with denser perturbations, even their drop in the strongest perturbation level is lower than that of human-written dialect prompts. This means that even when these LLMs are more familiar with AAVE, they still cannot perform as well in this dialect. Conversely, we find that Phi-3-Mini-128K-Instruct has better performance in AAVE data compared to perturbed texts of similar perplexities. This discrepancy seems to suggest that the small-scale model might have a different behavior pattern compared to larger models in dialect robustness. We further find that the denser AAVE features are, the bigger the performance drop is. We report further details in Appendix A.8, where we gradually control and inject synthetic lexico-syntactic dialect features following Ziems et al. (2022).
Generally, the findings in this subsection suggest that (i) the unfamiliarity of LLMs to AAVE does not explain the whole picture of the performance drop, so naively increasing AAVE in the training data may not diminish the performance gap, and (ii) LLMs, especially at large scales, might be even more brittle when facing the language of real users than what has been suggested by the previous robustness literature based on typo-style prompts (Zhu et al., 2023b).
4.2 Rephrasing Prompts in Standardized English Does Not Fill the AAVE Gap
Since LLMs generally show superior performance in Standardized ReDial, we experiment with instructing models to standardize and then answer the question to mitigate the AAVE bias, which we refer to as standardization. Specifically, we suffix ‘Let’s rephrase the query in Standard English first, then answer the question’ to every query. Results are reported in Figure 4 (bar plot).
Figure 4: Model pass rate and average response token count before and after being prompted for standardization. Standardization prompting generally improves LLM performance in both Standardized and AAVE ReDial (bar plot). However, even AAVE ReDial with standardization prompting cannot reach LLMs’ vanilla performance in Standardized ReDial, despite that they also tend to result in more tokens generated (scatter plot).
Indeed, LLM performance generally increases with standardization. Surprisingly, standardization improves model performance even when the prompt input is already in Standard English. Despite this, their performance on AAVE ReDial with standardization promoting still cannot reach their vanilla performance on Standardized ReDial. We further analyze the error patterns in Section 4.3.
Moreover, we notice that standardization introduces a computational overhead in terms of token count of LLMs’ responses (Figure 4, scatter plot), especially in GPT-4o and GPT-4. This means that even if dialect users pay more, they might still not be able to receive the same quality service as users who use Standardized English.
4.3 Qualitative Analysis
Intuitively, standardization prompting should cancel the dialect gap. However, we still observe a sensible gap between model performance on Standardized ReDial with zero-shot prompting and AAVE ReDial with standardization prompting. In this section, we qualitatively compare GPT-4o’s outputs in these two settings, the model with the best absolute overall performance, and examine its errors. We focus on the math subset of ReDial and identify three key error patterns: wrong question rephrasing, distraction by irrelevant information, and failure to execute all steps.
5 Related works

### v2

@@ 3.1.1 Models
We test five families of models, two proprietary and three open-source. The rationale is to benchmark widely used LLMs with impressive reasoning performance. All experiments were conducted between September and December 2024.
@@
We use GPT-o1 (OpenAI, ), GPT-4o (OpenAI, 2024), GPT-4 (Achiam et al., 2023), GPT-3.5-turbo (Achiam et al., 2023), [^4] Proprietary model API version information: o1: GPT-o1-preview; gpt-4o: 2024-05-13; gpt-4: 2024-05-03; gpt-3.5-turbo: 2023-11-06. as a family of closed-source models to compare with open-source models for dialect robustness. In particular, o1 is trained using large-scale reinforcement learning (RL) to reason through a chain of thoughts and scales inference time computation to achieve highly complex reasoning paths. It has demonstrated significant improvements in reasoning tasks (OpenAI, ). We use GPT-o1 model to understand how RL reasoning post-training affects LLMs’ dialect robustness and fairness.
Claude.
Developed by Anthropic, the Claude 3 model family represents a widely-used proprietary LLM. For our experiments, we utilize the Claude 3.5 Sonnet model  (Anthropic, ).
@@
We use Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). Mistral-7B-Instruct-v0.3 is reported to be outstanding in reasoning; with Mixtral-8x7B-Instruct-v0.1, we can understand whether Mixture-of-Expert (MoE) architectures enhance dialect robustness.
@@
We use Phi-3-Mini / Small / Medium-128K-Instruct (Abdin et al., 2024; Gunasekar et al., 2023) in our experiment. Phi-3 models, pre-trained on carefully designed “textbook” data, are reported for impressive performance in reasoning despite their small sizes (3.8/7/14B parameters each). We use these models to understand how highly curated pre-training data affect LLMs’ dialect robustness and fairness.
Model Setting Algorithm Logic Math Integrated All
SE AAVE SE AAVE SE AAVE SE AAVE SE AAVE $\Delta$
GPT-o1 Direct 0.818 0.825 0.947 0.923 0.878 0.815 0.942 0.925 0.892 0.866 -0.026
GPT-4o Direct 0.790 0.761 0.933 0.930 0.818 0.768 0.783 0.312 0.832 0.716 -0.116
CoT 0.771 0.761 0.950 0.920 0.815 0.771 0.762 0.662 0.826 0.784 -0.043
GPT-4 Direct 0.742 0.723 0.840 0.713 0.796 0.749 0.217 0.133 0.678 0.612 -0.067
CoT 0.723 0.608 0.920 0.813 0.793 0.743 0.283 0.058 0.706 0.590 -0.115
GPT-3.5-turbo Direct 0.653 0.631 0.667 0.443 0.533 0.544 0.200 0.129 0.531 0.460 -0.072
CoT 0.646 0.551 0.753 0.543 0.503 0.425 0.075 0.067 0.517 0.416 -0.101
Claude-3.5-Sonnet Direct 0.771 0.806 0.970 0.930 0.851 0.776 0.879 0.717 0.865 0.810 -0.055
CoT 0.774 0.736 0.953 0.940 0.859 0.796 0.900 0.771 0.868 0.811 -0.058
LLaMA-3.1-70B Direct 0.726 0.653 0.767 0.893 0.702 0.630 0.392 0.112 0.663 0.599 -0.064
CoT 0.723 0.653 0.880 0.870 0.809 0.768 0.579 0.500 0.759 0.711 -0.049
LLaMA-3-70B Direct 0.682 0.643 0.907 0.887 0.663 0.552 0.158 0.067 0.628 0.562 -0.066
CoT 0.697 0.646 0.923 0.887 0.616 0.561 0.517 0.350 0.693 0.622 -0.072
LLaMA-3-8B Direct 0.535 0.510 0.827 0.800 0.478 0.464 0.025 0.067 0.489 0.480 -0.009
CoT 0.532 0.478 0.827 0.800 0.475 0.492 0.029 0.025 0.488 0.472 -0.016
Mixtral-8x7B Direct 0.452 0.401 0.520 0.340 0.414 0.240 0.100 0.075 0.388 0.274 -0.114
CoT 0.468 0.411 0.687 0.567 0.384 0.285 0.133 0.071 0.431 0.345 -0.086
Mistral-7B Direct 0.331 0.255 0.400 0.213 0.315 0.271 0.096 0.075 0.297 0.214 -0.083
CoT 0.312 0.245 0.453 0.347 0.323 0.293 0.083 0.083 0.305 0.252 -0.053
Phi-3-Medium Direct 0.545 0.433 0.867 0.790 0.500 0.470 0.050 0.038 0.513 0.454 -0.059
CoT 0.548 0.455 0.860 0.827 0.492 0.439 0.067 0.029 0.513 0.458 -0.055
Phi-3-Small Direct 0.615 0.252 0.820 0.760 0.530 0.525 0.058 0.062 0.530 0.421 -0.109
CoT 0.570 0.194 0.893 0.843 0.544 0.522 0.096 0.079 0.549 0.429 -0.119
Phi-3-Mini Direct 0.557 0.427 0.520 0.550 0.605 0.525 0.021 0.042 0.456 0.410 -0.046
CoT 0.576 0.443 0.773 0.750 0.622 0.528 0.017 0.021 0.528 0.461 -0.067
Table 2: We report model pass rates using direct and CoT prompting on ReDial, including individual performances on subtasks and overall performance/gap (in column all). We follow the recommendations from Dror et al. (2018) and test the statistical significance of performance differences between SE and AAVE on all results using the McNemar’s test for binary data (McNemar, 1947). We correct p-values for multiple measurements using the Holm-Bonferroni method (Holm, 1979). Statistically significant drops are in bold. Details in Appendix A.7.
@@ 3.1.2 Implementation and Evaluation
We set the temperature to zero for the main experiments to ensure maximum reproducibility. We report two prompting methods in our main results: (i) direct prompting LLMs with task instances, which resembles general real-life use cases the most (Direct) and (ii) zero-shot Chain of Thought (Wei et al., 2022; Kojima et al., 2022) (CoT, i.e., adding instructions in the spirit of “Let’s think step by step” on top of task descriptions, which resembles expert user prompts to improve model performance). For GPT-o1, we only test the direct prompting as its inherent CoT reasoning pattern. [^5] We also test non-zero temperatures and report results in Appendix A.6. We report further implementation details in Appendix A.5.666We deliberately avoid testing advanced prompting methods, such as Tree of Thought (Yao et al., 2024) and Self-Refine (Madaan et al., 2024). Our focus is on evaluating how LLMs perform when prompted for everyday use by dialect users, which is critical for assessing fairness in LLMs. Similarly, we do not fine-tune any models, as our study aims to investigate biases inherent in base models. The effects of fine-tuning are beyond the scope of this study.
Algorithm Math Logic Intergrated All
SE 0.632 0.622 0.768 0.302 0.597
AAVE 0.563 0.564 0.706 0.212 0.529
$\Delta$0.069 -0.058 -0.062 -0.090 -0.068
Table 3: Pass rates by task averaged across responses from all models with direct prompting. In bold, results statistically deviate according to McNemar’s tests applied to AAVE and SE (i.e., models have significant drops in AAVE). We also report the SAE-AAVE absolute delta in performance.
@@ 3.2 Experimental Results
We report pass rates for ReDial in Table 2 and  3, and summarize the main results of our experiments.
All Models are Brittle.
@@
All models experience performance drops in AAVE compared to SE ReDial, and these drops are statistically significant in all cases, with the sole exception of LLaMA-3-8B-Instruct. This indicates that our benchmark poses huge challenges to models, both in terms of absolute performance and with respect to their dialect robustness and fairness.
The absolute performance gaps commonly range from around 5% to over 10% ($\Delta$ in table 2). Specifically, GPT-4o (zero-shot) shows an absolute gap of 11.6%, dropping from an average of 0.832 to 0.716. GPT-4 (CoT) exhibits an 11.5% drop. Mixtral-8x7B-Instruct-v0.1 (zero-shot) shows a particularly large difference of 11.4% points as well. Interestingly, we found that although the performance drop of GPT-o1 is smaller than other GPT models, but still significant. It indicates that although further RL post-training on general reasoning and inference scaling can systematically enhance dialect robustness and fairness, they cannot completely solve the problem.
In short, dialect unfairness and brittleness are identified in all the models we examined, including the MoE and large reasoning models. This finding indicates that the problem is widespread, non-trivial, and cannot be easily mitigated by naively changing model architecture or proposing more complex reasoning paths.
All Tasks are Brittle.
When aggregated by task type, AAVE queries cause a statistically significant performance drop across all these categories (Table 3). For instance, when averaging results across all models: direct prompting leads to an average 10% relative performance drop.
Interestingly, integrated reasoning tasks, which require multiple reasoning skill compositions, show some of the largest relative drops (about 30%). This suggests that compositionally complex task may be more prone to dialect brittleness.
Prompting and Inference Scaling are Brittle.
While CoT prompting can slightly reduce the discrepancy for some models, it fails to close it entirely. For example, GPT-4o’s performance gap decreases from about 0.116 (zero-shot) to 0.055 (CoT). This suggests that even when models are given additional reasoning "scaffolding," their understanding and performance in AAVE remain comparatively weaker than in SE, which is also in line with our observation with GPT-o1 results.
Model Scaling is Brittle.
All model families display some degree of dialect-related performance degradation. A notable observation is that simply using larger models does not inherently improve robustness to AAVE. For example, even LLaMA-3.1-70B-Instruct, among the largest and most capable tested models, suffers from significant performance drops on AAVE queries. This pattern holds across the board, indicating that scaling model size alone is insufficient to address dialect-related performance disparities.
4 Discussions
This section investigates the potential reasons for AAVE’s brittleness. We show that LLMs’ brittleness with AAVE reasoning queries cannot be simply attributed to the lack of understanding of this dialect or simple lexical features. The nuanced conversational norms of AAVE also play an important role in the problems for LLMs.
4.1 General Understanding and Morphosyntactic Features
One possible explanation for the performance drop is that LLMs cannot process AAVE dialect. We thus computed perplexities on ReDial AAVE vs. SAE prompts. Indeed, Table 4 confirms that LLMs exhibit higher perplexities on dialect than SE.
Models SE AAVE
LLaMA-3.1-70B-Instruct 9.4 17.5
Phi-3-Medium-128K-Instruct 5.9 12.5
Phi-3-Mini-128K-Instruct 7.1 15.9
@@ 4.1 General Understanding and Morphosyntactic Features
Table 4: Averaged perplexities across instances calculated by different models on SE/AAVE ReDial.
However, is the insufficient understanding the only reason that is causing LLMs’ performance to drop? To answer it, we gradually inject typos in SE ReDial by replacing/deleting/adding words/characters, such that we make the input texts more difficult for LLMs (i.e., the measured perplexity goes up). We find that while these perturbations degrade model performance, the drop does not reach the severity observed with natural AAVE data on large-scale models (see full results in Section A.8). This discrepancy suggests that AAVE brittleness is not solely due to the general difficulty in processing. There are some dialect-specific reasons for the performance drop.
If language-agnostic processing ability cannot explain LLMs’ brittleness, can we attribute the problem to morphosyntactic AAVE features? Following Ziems et al. (2022, 2023), we use morphosyntactic transformation rules to inject AAVE features into SE ReDial. We find that performance degradation generally intensifies as the density of AAVE-specific features increases when we gradually inject AAVE features into standard ReDial (see full results in Section A.9). This suggests that these features play a significant role in diminishing model performance.
However, even under the most extreme synthetic perturbations, performance drops are notably less severe than those observed with human-rewritten prompts. This underscores the critical importance of our high-quality human-annotated dialect data ReDial for evaluating LLM fairness and robustness. Synthetic rule-based transformations provide valuable insights, yet fail to capture the contextual depth of real-world dialect usage.
4.2 AAVE Conversational Norms
We use the mutual information between the token distributions of SE and AAVE ReDial to find that the top 5 most informative AAVE features in terms of distinguishing them from SE are , up, in, gon, and gotta. We note that many of these features are not well-known AAVE-specific features (e.g., up). Through a further investigation of our dataset, we find that these lexicons are associated with phrase-level AAVE constructions. For instance, instead of saying …encode the answer… in Standard English, AAVE instruction says …wrap it up…. This finding is particularly interesting as we find that, in addition to previous linguistic observations of AAVE-specific morphosyntactic features, there are conversational norms of the dialect such as nuanced uses of phrases.
We compute Spearman’s correlation between the frequency of the features we find in each instance and their corresponding performance drop. Indeed, these features play a significant role in predicting GPT-4o’s performance degradation (r=-0.318, p<0.001). We further implement and analyze 12 rule-based AAVE features following Ziems et al. (2022) (details in Appendix A.11), which are well documented in linguistic literature such as "finna" as a maker of immediate future Nguyen and Grieve (2020). We notice that the influential lexical features are a subset of the feature set discovered by mutual information (i.e., some of the actual influential features are not encoded in synthetic transformation rules). Consequently, the influence of synthetic features is not as strong as those discovered by mutual information (r=-0.256, p<0.001). This means that simple rule-based transformations that implement the most salient morphosyntactic AAVE features may not be able to capture rich, context-dependent use of the dialect and, therefore fall short in predicting LLMs’ performance in real workflows.
Aided by GPT-o1 preview to filter the vast amount of data, we conducted a linguistically informed analysis of LLMs’ most frequent errors on AAVE. For the algorithmic tasks, grammatical constructions and non-standard verb forms (e.g., finna’, ’em), omission of articles and auxiliary verbs may cause the model to misinterpret references and function naming conventions. For example, GPT-4o interprets you gon’ write a python function, python_function as a general statement rather than a directive to name the function. On logic tasks, the frequent use of double negatives, zero copula, and inverted conditionals introduces structural ambiguities. For example, the construct He don’t take no breaks can invert intended meanings, leading the model to misunderstand conditional statements. On math, informal expressions, unclear quantity references, and non-standard comparative cause erroneous parsing of numerical information and confusion over collective versus individual quantities. Informal phrasings like half as much as he be runnin’ and ambiguous comparative expressions (4 fewer boxes of apple pie than on Sunday) can cause the model to misinterpret numerical relationships, resulting in errors when calculating totals, differences, or fractions. On the integrated task, phonetic spellings, colloquial connectors, and inverted word orders limit the model’s understanding of concurrency, and stepwise instructions. Such dialectal nuances highlight the necessity of our dataset and also call for more efforts to collect more human data for relevant purposes.
5 Related Work